The number of policies for 5 equal partners in a cross-purchase arrangement is:

A. 5

B. 10

C. 15

D. 20
Answer #1

Correct answer is D.20

No. of polices in cross purchase arrangement = n*(n-1)

= 5*(5-1)

= 5*4

= 20

Please Thumbs Up, Thank you.